OpenStudy (anonymous):

What is the slope-intercept form of the function described by this table? x 1 2 3 4 y 1 −2 −5 −8 y= ___x + ___

OpenStudy (susanna):

You can graph it, and see from the graph; or you can do what we did before. Take any 2 points (1,1)(2, -2) find the slope plug in 1 of the points and the slope into y-y=m(x-x) and solve for y.

OpenStudy (susanna):

\[\frac{ -2-1 }{ 2-1}=\frac{ -3 }{ 1 }=-3\]

OpenStudy (susanna):

that's the slope

OpenStudy (anonymous):

alright good.

OpenStudy (susanna):

now, take any point (1, 1) y-1=-3(x-1)

OpenStudy (tkhunny):

You probably should verify that it is also the slope between other pairs of points. Alternatively, demonstrating the arithmetic progression of the y-values might be a useful exercise.

OpenStudy (anonymous):

3x - 2?

OpenStudy (anonymous):

hm? what do u mean?

OpenStudy (susanna):

y-1=-3x+3

OpenStudy (susanna):

add 1 to both sides

OpenStudy (anonymous):

3x + 4

OpenStudy (anonymous):

-3*

OpenStudy (susanna):

\[\frac{ -5+8 }{ 3-4 }=\frac{ 3 }{ -1 }=-3\] @tkhunny you are correct, y=-3x+4